# A Robust Autoencoder Ensemble-Based Approach for Anomaly Detection in Text

### v1

@@ Abstract
In this work, a robust autoencoder ensemble-based approach designed to address anomaly detection in text corpora is introduced. Each autoencoder within the ensemble incorporates a local robust subspace recovery projection of the original data in its encoding embedding, leveraging the geometric properties of the k-nearest neighbors to optimize subspace recovery and identify anomalous patterns in textual data. The evaluation of such an approach needs an experimental setting dedicated to the context of textual anomaly detection. Thus, beforehand, a comprehensive real-world taxonomy is introduced to distinguish between independent anomalies and contextual anomalies. Such a study to identify clearly the kinds of anomalies appearing in a textual context aims at addressing a critical gap in the existing literature. Then, extensive experiments on classical text corpora have been conducted and their results are presented that highlights the efficiency, both in robustness and in performance, of the robust autoencoder ensemble-based approach when detecting both independent and contextual anomalies. Diverse range of tasks, including classification, sentiment analysis, and spam detection, across eight different corpora, have been studied in these experiments.
@@ 1 Introduction
Anomaly detection Chandola et al. (2009); Ruff et al. (2021) and outlier detection Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017a) are tasks aiming to estimate whether an observation is normal or not. Both of these tasks are referring to similar characteristics, and the terms anomaly and outlier are interchangeable in this paper. Several works have been devoted to study and categorise the different characteristics of an anomaly, and three principal kinds of anomalies can be noted: independent, contextual and collective. Defining a taxonomy for anomalies is crucial while the task to find anomalies lacks dedicated datasets. Such issue often leads practitioners to introduce an artificial contamination to produce anomalies into datasets. With the emergence of new machine learning methods and the availability of many datasets and corpora, anomaly detection can be addressed through many approaches Beggel et al. (2020); He et al. (2021); Goodge et al. (2023), however, in most cases, models are based on one-class classification (OCC) Khan and Madden (2014); Ruff et al. (2018); Sohn et al. (2021).
@@
Performing anomaly detection on textual data is less common than on many other types of data (images, time series and medical) but it comes with several useful applications that helps discerning wrong web content, hateful messages, spam or also errors in news feed. One of the main difficulties in this domain is the reproduction of experimental protocols and results from the literature Bejan et al. (2023). Indeed, there is a great difference between tackling independent outliers and contextual outliers Mahapatra et al. (2012); Fouché et al. (2020); Bejan et al. (2023) using semantic in text. For the former, anomaly classifier needs to differentiate two kinds of documents that come from unrelated topics (eg. sports and computer) but for the latter, they needs to detect a topic contaminated with another "sibling" (see Fig. 1 for examples of this difference). It appears that most of the recent works contaminate corpora without addressing the problem of which kind of anomaly/outlier is added in this process Manevitz and Yousef (2001); Kannan et al. (2017); Ruff et al. (2019a); Lai et al. (2020).
In this paper, based on recent literature, we address the problem of taxonomy after defining two kinds of anomaly: independent and contextual. Those anomalies are dependent of any semantic hierarchy and we propose in the experiments section to detail furthermore how they are benchmarked. For tackling both kind of anomalies, we propose a novel ensemble-based approach built with a robust autoencoder.
This paper is composed as follows. In Section 2, recent works are presented. In Section 3, our approach to build an Ensemble robust subspace local recovery encoder is introduced with a description of its properties. Several experiments are presented in Section 4. Finally, Section 6 concludes and presents some future work.
@@ 2 Related works
Recent advances in word embedding with language models like BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019), and large language models such as Llama Touvron et al. (2023) and OPT Zhang et al. (2022) have shown promising characteristics for anomaly detection. However, only few methods of the literature propose their use Manolache et al. (2021); Ruff et al. (2019a). Other methods like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Textual Outlier using Nonnegative Matrix Factorization (TONMF) Kannan et al. (2017) rely on a TF-IDF representation. Surprinsinly, recent methods are not using outlier ensemble methods Aggarwal and Sathe (2015); Zhao et al. (2019a); Zimek et al. (2014) for performing anomaly/outlier detection with text data. Their usage in both Mixture of Experts Du et al. (2022) and ensemble approach Briskilal and Subalalitha (2022) in various NLP tasks highlights that anomaly detection in text data misses important approaches.
@@
AutoEncoders (AE) have been widely used for anomaly/outlier detection with high-dimensional data Chen et al. (2017); Kieu et al. (2019) and are also successful with other kind of data An and Cho (2015); Chen et al. (2018); Lai et al. (2020); Zhou and Paffenroth (2017); Beggel et al. (2020). The risk of using autoencoders with language models leads to the apparition of degenerate solution in the learning step. Robust properties are needed in such scenario for handling the manifold collapse phenomenon Wang and Wang (2019). Such issue can be mitigated using robust projections in the learned space of the autoencoder (eigenfunction decomposition Bengio et al. (2004), …). The Robust Subspace Recovery Lerman and Maunu (2018); Rahmani and Atia (2017) is a robust manifold learning technique that map inlier distributions in a subspace where anomalies/outliers are at the edge. An autoencoder based on such an approach Lai et al. (2020) encodes observation from the input space in a robust subspace. One problem of such autoencoder is the difficulty to address scenarios with features that present similar range values (problem of entanglement). This entanglement is especially more difficult with contextual anomalies where documents shares similar semantic properties. Such problem can be tackled using a local projection from the nearest neighbours in the latent representation space. Several manifold learning approaches like locally linear embedding (LLE) Roweis and Saul (2000), neighbourhood components analysis (NCA) Goldberger et al. (2004) or soft nearest neighbours (SNN) Frosst et al. (2019) can help to mitigate it.
In this paper we introduce an ensemble-based approach with randomly connected autoencoders. We also present a new autoencoder, robust subspace local recovery autoencoder (RLAE), that performs robust subspace representation with local neighbouring recovery. In this approach, called Ensemble robust subspace Local recovery Autoencoder (ErLA), a RLAE ensemble, it can be seen as an ensemble of several subspace that aims to find normal data with different manifolds. It is based on the hypothesis that one distribution is highly contaminated, and that inliers lies in a low-dimensional subspace.
3 ErLA: Ensemble robust subspace Local recovery Autoencoder
@@
This section presents our approach ErLA and a description of its properties. While robust subspace recovery autoencoders have successfully tackled anomaly detection in text corpora, they lack locality and geometry awareness for mitigating disentanglement and manifold collapse in transformer-based language models. Thus, we introduce RLAE which integrates locality in the latent representation through a local neighbouring method.
The section is structured with a presentation of the randomly connected autoencoders, followed by a presentation of RSR loss. We then introduce the locally linear embedding loss term of ErLA before presenting its ensemble method. Finally, we present the representation of text.
@@ 3.1 Randomly Connected One-Class Autoencoder
Instead of using fully connected autoencoders, we propose to use randomly connected autoencoders. In the case of RSRAE, it is a novel approach and allow us to build ensemble autoencoders with different base detectors.
Let $\mathrm{X}$ be a dataset of $\mathrm{N}$ instances such as $\mathrm{X}=\{\mathrm{x_{1}},...,\mathrm{x_{N}}\}$. Each instance has $\mathrm{D}$ dimensions which correspond to its attributes: $\mathrm{x_{i}}=\{x_{1},...,x_{D}\}$. An Autoencoder is a neural networks in which the encoder $\mathcal{E}$ maps an instance $\mathrm{x_{i}}$ in a latent representation noted $\mathrm{z}_{i}=\mathcal{E}(\mathrm{x}_{i})\in\mathbb{R}^{e}$ of dimension $e$. The RSR layer is a linear transformation $\mathbf{A}\in\mathbb{R}^{d\times e}$ that reduces the dimension to $d$. We denote $\mathrm{\hat{z}_{i}}$ the representation of $\mathrm{z_{i}}$ through the RSR layer, such as $\mathrm{\hat{z}_{i}}=\mathbf{A}\mathrm{z_{i}}\in\mathbb{R}^{d}$. The decoder $\mathcal{D}$ maps $\mathrm{\hat{z}_{i}}$ to $\mathrm{\hat{x}_{i}}$ in the original space $D$. The matrix $\mathbf{A}$ and the parameters of $\mathcal{E}$ and $\mathcal{D}$ are obtained with the minimization of a loss function.
Similarly to Chen et al. (2017) we introduce autoencoders with random connection such as we increase the variance of our model. In the autoencoders ensemble each autoencoder has a random probability of having several of its connections to be cut. Thus, we set the probability of disconnection with a random rate in $[0.2,0.5]$.
@@ 3.2 Robust Subspace Recovery Layer
We present the RSR AutoEncoder that aims to robustly and nonlinearly reduce the dimension of the original data Lerman and Maunu (2018). The RSR layer maps the inliers around their original locations and the outliers far from their original locations. The loss function minimises the sum of the autoencoder loss function noted $L_{AE}$ with the RSR loss function noted $L_{RSR}$.
@@
For performing the subspace recovery, we denote two terms that have different roles in the minimisation process. The first term enforces the RSR layer to be robust (PCA estimation) and the second enforces the projection to be orthogonal:
@@ 3.3 Robust local embedding
Locally Linear Embedding (LLE) Roweis and Saul (2000); Chen and Liu (2011) is a popular dimensionality reduction technique that aims to preserve the local geometry of the data in a lower-dimensional subspace. It is based on the assumption that data points in a local neighbourhood can be linearly represented by their neighbouring data points. The LNE term in the loss function encourages the autoencoder to learn representations that preserve the relationships between data points in their local neighbourhoods. By doing so, it helps to project the Euclidean distance with its neighbours in the learned subspace. Based on Equation 2, the reconstruction loss function of RSRAE enforces robustness with $L^{1}_{AE}$ and the orthogonality with $L^{1}_{RSR}$. Because the learned representation of the encoder is compressed in a $e$ dimension space, the locality of the subspace is not handled.
For tackling this problem, we propose to introduce a third term to $L_{RLAE}$ based on a local neighbours embedding (LNE). Given a set of data points $\{\mathrm{x_{i}}\}_{i=1}^{N}$ in the input space, the goal of LNE is to find a lower-dimensional representation $\{\mathrm{z_{i}}\}_{i=1}^{N}$ in the output space (the subspace learned by the autoencoder) such that the local relationships between data points are preserved. We note:
@@
The LNE term encourages the autoencoder to find a representation for each data point as a combination of its k-nearest neighbours in the input space. With minimisation of the LNE term in the loss function, the autoencoder learns to preserve the local linear relationships, which ultimately helps to project the Euclidean distance with its neighbours. The reconstruction errors of LNE is measured by the cost function:
@@
The weight $w_{j}$ assigned to the neighbour $\mathrm{x_{ij}}$ in the local reconstruction of $\mathrm{x_{i}}$ is determined based on its corresponding distances. The inclusion of the LNE term in the loss function encourages the autoencoder to preserve the local geometric structure of the data in the learned subspace.
Finally, the RLAE cost function is measured as follows:
@@ 3.4 Ensemble Learning
The main idea behind ensemble methods is that a combination of several models, also called base detectors, and their outputs is more robust than usage of a single model. Such robustness can be observed against the bias-variance tradeoff and also for tackling the issue of overfitting.
Although the possibility to combine multiple base detectors is intuitive, the design of such approaches needs special attention regarding normalisation of outputs. In ErLA, we use the reconstruction error of each autoencoders and then we normalise each base detector scores through the standard deviation of one unit. We then take the median value for each observation.
@@ 3.5 Text Representation
Recent works Ruff et al. (2019a); Manolache et al. (2021) have recorded their results on BERT-based language models. Thus, we use RoBERTa Liu et al. (2019) for text representation. The ErLA model is not based on the self-attention mechanism, such as for Ruff et al. (2019a); Manolache et al. (2021), and we use the implementation of Reimers and Gurevych (2019).
4 Experiments
In this section we present several experiments on both independent anomalies and contextual anomalies. We present all corpora and how we perform this study. The section is structured as follows: i) presentation of the experimental protocol for tackling independent and contextual anomalies; ii) presentation of all corpora and their principal characteristics; iii) complete description of experiment settings; iv) results of conducted benchmarks.
Algorithm 1 TAC: Textual Anomaly Contamination
@@ 4 Experiments
4.1 Experimental protocol
A large number of contributions have settled anomaly/outlier taxonomies Hawkins (1980); Hodge and Austin (2004); Zhang (2013); Aggarwal (2017b); Ruff et al. (2019a) several types of outliers have been proposed in the literature: Point anomaly, Conditional/Contextual anomaly and Collective/Group anomaly. A similar taxonomy can be applied to textual data. Consequently, various types of outliers frequently coexist within the documents of a given corpus. The definition of a topic can be assimilated to the subject matter that a document addresses. Depending on the document type, there may be multiple subtopics within a broader category (e.g., a sports topic that encompasses football and tennis). Thus, accounting for this hierarchical structure introduces a form of contextual anomaly. These contextual outliers may appear unremarkable in isolation but are considered outliers when associated with a small subset of the corpus.
Collective anomaly poses challenges in terms of formalisation due to the contextual nature of textual data. To illustrate this, consider a legal document mentioning a football player, which would be anomalous if incorrectly appearing in a sports-related corpus. Point outliers represent observations that lack any meaningful relationship with other topics. Specifically, anomaly topics and inlier topics have different hierarchical parents within the category structure. Let a labelled document of a corpus $(\mathrm{x},y)\in\mathrm{X}\times\mathcal{Y}$ and $\zeta$ be the inlier category, and its corresponding subset $\mathrm{X}_{\zeta}\subseteq\mathrm{X}$. We define $\mathcal{A}$ the subset of all anomalies such as $\mathcal{A}\subset\mathrm{X}$. We have:

### v2

@@ Abstract
Anomaly detection (AD) is a fast growing and popular domain among established applications like vision and time series. We observe a rich literature for these applications, but anomaly detection in text is only starting to blossom. Recently, self-supervised methods with self-attention mechanism have been the most popular choice. While recent works have proposed a working ground for building and benchmarking state of the art approaches, we propose two principal contributions in this paper: contextual anomaly contamination and a novel ensemble-based approach. Our method, Textual Anomaly Contamination (TAC), allows to contaminate inlier classes with either independent or contextual anomalies. In the literature, it appears that this distinction is not performed. For finding contextual anomalies, we propose RoSAE, a Robust Subspace Local Recovery Autoencoder Ensemble. All autoencoders of the ensemble present a different latent representation through local manifold learning. Benchmark shows that our approach outperforms recent works on both independent and contextual anomalies, while being more robust. We also provide 8 dataset comparison instead of only relying to Reuters and 20 Newsgroups corpora.
@@ 1 Introduction
Anomaly Detection (AD) Chandola et al. (2009); Ruff et al. (2021) and Outlier Detection (OD) Hodge and Austin (2004); Zhang (2013); Aggarwal (2017a) are tasks aiming to estimate whether an observation is normal or not. These tasks have been recently applied on text for highlighting particular or abnormal texts in a corpus. In this context, it is still a blossoming and promising field. Nowadays, most popular approaches are performing One-Class Classification (OCC) for discerning outliers (anomalies) from inliers Kannan et al. (2017); Lai et al. (2020); Fouché et al. (2020), and self-supervised approaches are often preferred for text Ruff et al. (2019); Manolache et al. (2021); Das et al. (2023). Self-supervised approaches are not limited to text, and applications with large language model for vision exists Wang et al. (2019); Zhou et al. (2023).
@@
Ensemble-based approaches are not present in the literature, despite their great success on various kind of data Chen et al. (2017); Briskilal and Subalalitha (2022). Whereas those approaches benefits of decision robustness and reduced variance. Most of recent approaches are using Multi-Head Self-Attention Ruff et al. (2019); Manolache et al. (2021); Das et al. (2023), learning a representation from pre-trained language models as BERT-based models Devlin et al. (2019); Liu et al. (2019); Reimers and Gurevych (2019). An underlying problem of AD lies in the identification of different kinds of anomalies. Recent approaches does not usually consider these differences when building a detection model. While independent, contextual and collective anomalies have been identified for other kind of data Chandola et al. (2009); Aggarwal (2017b), it is different for text. Usually text anomalies are only considered as documents from one topic appearing in the wrong topic subset (Figure 1). However anomalies can have multiple forms, and independent anomalies are easier to find than contextual anomalies.
In this paper we introduce a novel ensemble-based approach that performs contextual anomaly detection. In comparison to state-of-the-art approaches, we focus first on the identification of two kinds of anomaly, independent and contextual, and introduce the Textual Anomaly Contamination (TAC), a novel experimental setup. Also, our main contribution lies in the introduction of a novel approach for detecting anomalies which is based on a robust autoencoder ensemble. Our autoencoder, Robust Subspace Local Recovery AutoEncoder (RLAE), faces contaminated corpora (trained on both inliers and anomalies) and discriminates inliers from anomalies under the assumption that one distribution is highly contaminated. Thus, our Robust Subspace Local Recovery Autoencoder Ensemble (RoSAE) represents inliers in a low-dimensional subspace and proposes different manifold projections for one observation.
@@ 2 Related works
Recent advances in word embedding with language models like BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019), and large language models such as Llama Dubey et al. (2024) and OPT Zhang et al. (2022) have shown promising characteristics for anomaly detection. However, only few methods of the literature propose their use Ruff et al. (2019); Manolache et al. (2021); Das et al. (2023). Other methods like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Textual Outlier using Nonnegative Matrix Factorization (TONMF) Kannan et al. (2017) rely on a TF-IDF representation. Surprisingly, recent methods are not using outlier ensemble methods Aggarwal and Sathe (2015); Zhao et al. (2019a); Zimek et al. (2014) for performing anomaly/outlier detection with text data.
@@
AutoEncoders (AE) have been widely used for anomaly/outlier detection with high-dimensional data Chen et al. (2017); Kieu et al. (2019) and are also successful with other kind of data An and Cho (2015); Chen et al. (2018); Lai et al. (2020); Zhou and Paffenroth (2017); Beggel et al. (2020). The risk of using autoencoders with language models leads to the apparition of degenerate solution in the learning step. Robust properties are needed in such scenario for handling the manifold collapse phenomenon Wang and Wang (2019). Such issue can be mitigated using robust projections in the learned space of the autoencoder (eigenfunction decomposition Bengio et al. (2004), …). The Robust Subspace Recovery Lerman and Maunu (2018); Rahmani and Atia (2017) is a robust manifold learning technique that map inlier distributions in a subspace where anomalies/outliers are at the edge. Based on such an approach, an autoencoder Lai et al. (2020) encodes observation from the input space in a robust subspace. One problem of such autoencoder is the difficulty to address scenarios with features that present similar range values (problem of entanglement) Wang et al. (2022). Entanglement is especially more difficult with contextual anomalies where documents shares similar semantic properties. Such problem can be tackled using a local projection from the nearest neighbours in the latent representation space. Several manifold learning approaches like locally linear embedding (LLE) Roweis and Saul (2000), neighbourhood components analysis (NCA) Goldberger et al. (2004) or soft nearest neighbours (SNN) Frosst et al. (2019) can help to mitigate it.
Figure 2: Overview of our global system, including contamination setup and our proposed approach RoSAE.
A large number of contributions have settled anomaly/outlier taxonomies Zhang (2013); Aggarwal (2017b); Ruff et al. (2019) and several types of anomalies have been proposed in the literature: Independent anomaly, Contextual anomaly and Collective anomaly. Consequently, various types of anomalies frequently coexist within one corpus. A topic is defined as the main subject of a document Bejan et al. (2023). Depending on the document, there may be multiple subtopics within a broader category (e.g., a sports topic that encompasses football and tennis). Thus, accounting for this hierarchical structure, a form of contextual anomaly arises Fouché et al. (2020).
Three different approaches have been used for performing corpus contamination. The first is a manually picked contamination Manevitz and Yousef (2001); Kannan et al. (2017), using text classification corpora like 20 Newsgroups, where topics are unrelated. A more practical contamination process is performed by recent works like Lai et al. (2020); Fouché et al. (2020), in which the inlier topic is selected and all other topics are considered anomalies. In this scenario, the hierarchical structure of topics is not handled, and related topics are mixed with each others. Finally, more recently, corpora have been prepared with careful consideration of topic hierarchy Ruff et al. (2019); Manolache et al. (2021); Das et al. (2023). We also note that there exists different kind of training procedure. In Ruff et al. (2019); Manolache et al. (2021) the proposed model is trained on inlier class, and then tested on contaminated split. For Lai et al. (2020); Fouché et al. (2020); Das et al. (2023), the training and testing step is performed on contaminated split.
3 RoSAE: Robust Subspace Local Recovery Autoencoder Ensemble
@@
While robust subspace recovery autoencoders have successfully tackled anomaly detection in text corpora, they lack locality and geometry awareness for mitigating disentanglement and manifold collapse in transformer-based language models. Thus, we introduce Robust Subspace Local Recovery AutoEncoder (RLAE) which integrates locality in the latent representation through a local neighbouring method before presenting our ensemble-based approach RoSAE.
@@ 3.1 Randomly Connected One-Class Autoencoder
Let $\mathrm{X}$ be a dataset of $\mathrm{N}$ instances such as $\mathrm{X}=\{\mathrm{x_{1}},...,\mathrm{x_{N}}\}$. Each instance has $\mathrm{D}$ dimensions which correspond to its attributes: $\mathrm{x_{i}}=\{x_{1},...,x_{D}\}$. An Autoencoder is a neural networks in which the encoder $\mathcal{E}$ maps an instance $\mathrm{x_{i}}$ in a latent representation noted $\mathrm{z}_{i}=\mathcal{E}(\mathrm{x}_{i})\in\mathbb{R}^{e}$ of dimension $e$. The Robust Subspace Recovery (RSR) layer is a linear transformation $\mathbf{A}\in\mathbb{R}^{d\times e}$ that reduces the dimension to $d$. We denote $\mathrm{\hat{z}_{i}}$ the representation of $\mathrm{z_{i}}$ through the RSR layer, such as $\mathrm{\hat{z}_{i}}=\mathbf{A}\mathrm{z_{i}}\in\mathbb{R}^{d}$. The decoder $\mathcal{D}$ maps $\mathrm{\hat{z}_{i}}$ to $\mathrm{\hat{x}_{i}}$ in the original space $D$. The matrix $\mathbf{A}$ and the parameters of $\mathcal{E}$ and $\mathcal{D}$ are obtained with the minimization of a loss function.
Similarly to Chen et al. (2017) we introduce an autoencoder with random connection such as we increase the variance of our model. In the autoencoder ensembles, each autoencoder has a random probability of having several of its connections to be pruned. Thus, we set the probability of disconnection with a random rate in $[0.2,0.5]$.
@@ 3.2 Robust Subspace Recovery Layer
The RSR layer maps the inliers around their original locations, while outliers are projected far from their original locations Lerman and Maunu (2018); Lai et al. (2020). The loss function minimises the sum of the autoencoder loss function noted $L_{AE}$ with the RSR loss function noted $L_{RSR}$.
@@
For performing the subspace recovery, we denote two terms that have different roles in the minimisation process. Two terms are presented, with the first term enforcing the RSR layer to be robust (PCA estimation), and the second enforcing the projection to be orthogonal:
@@ 3.3 Robust Local Embedding
Locally Linear Embedding (LLE) Roweis and Saul (2000); Chen and Liu (2011) proposes the assumption that a local geometry can be linearly represented with surrounding observations. The Local Neighbouring Embedding (LNE) term in the loss function encourages the autoencoder to learn representations that preserve the relationships between data points and their local neighbours. Based on Equation 2, the reconstruction loss function enforces robustness with $L^{1}_{AE}$ and the orthogonality with $L^{1}_{RSR}$. Because the learned representation of the encoder is compressed in a $e$ dimension space, the locality of the subspace is not handled.
Based on LNE, we introduce a third term. Given a set of data points $\{\mathrm{x_{i}}\}_{i=1}^{N}$ in the input space, the goal of LNE is to find a lower-dimensional representation $\{\mathrm{z_{i}}\}_{i=1}^{N}$ in the output space (the subspace learned by the autoencoder) such that the local relationships between data points are preserved. We note:
@@
The LNE term encourages the autoencoder to find a representation for each data point as a combination of its k-nearest neighbours in the latent space. With minimisation of the LNE term in the loss function, the autoencoder learns to preserve the local linear relationships, which ultimately helps to project the Euclidean distance with its neighbours. The reconstruction errors of LNE is measured by the cost function:
@@
The weight $w_{j}$ assigned to the neighbour $\mathrm{x_{ij}}$ in the local reconstruction of $\mathrm{x_{i}}$ is determined based on its corresponding distances. The inclusion of the LNE term in the loss function encourages the autoencoder to preserve the local geometric structure of the data in the latent subspace.
Finally, the Robust Subspace Local Recovery AutoEncoder (RLAE) loss function is measured as follows:
@@ 3.4 Ensemble Learning
The main idea behind ensemble methods is that a combination of several models, also called base detectors, is more robust than usage of a single model. In RoSAE, we use the reconstruction error of each autoencoders and then we normalise each base detector scores through the standard deviation of one unit. We then take the median value for each observation.
@@ 3.5 Text Representation
Recent works have recorded their results on BERT-based language models. The RoSAE model is not based on the self-attention mechanism, such as for Ruff et al. (2019); Manolache et al. (2021), and we use the RoBERTa Sentence Transformer from Reimers and Gurevych (2019).
Algorithm 1 TAC: Textual Anomaly Contamination (Independent)
@@
4 Textual Anomaly Contamination (TAC)
Text AD is lacking dedicated dataset, resulting in all state-of-the-art methods preparing benchmark corpora with synthetic contamination. We propose a contamination algorithm for performing contamination with independent and contextual anomalies (Figure 2). To illustrate this, let a legal document mentioning a football player, it would be anomalous if incorrectly appearing in a sports-related corpus. Independent anomalies represent observations that lack any meaningful relationship with other topics. Specifically, anomaly topics and inlier topics have different hierarchical parents within the category structure. Let a corpus of text $\mathrm{X}$ with its labelled topic $\mathcal{Y}$, a labelled document of a corpus $(\mathrm{x},y)\in\mathrm{X}\times\mathcal{Y}$ and $\zeta$ the inlier category The corresponding inlier subset is noted $\mathrm{X}_{\zeta}\subseteq\mathrm{X}$. We define $\mathcal{A}$ the subset of all anomalies such as $\mathcal{A}\subset\mathrm{X}$. We have: